A small airport is served by four airlines: A, B, C which offer overseas flights, and D which offers domestic flights.

Over a one month period, the number of on-time and delayed arrivals are recorded for all four airlines. They are shown in the table below.

What is the probability that a flight was delayed and it was run by airline D? Give your answer correct to four decimal places.



Airline On-Time Delayed

A 105 15

B 38 20

C 21 7

D 200 59

1
Expert's answer
2021-02-24T06:58:00-0500

"\text{let }S -\text{number of all aircraft departures}"

"S = 105+15+38+20+21+7+200+59=465"

"\text{let }S_d -\text{number of all delayed aircraft departures}"

"S_d=15+20+7+59=101"

"\text{delayed departure probability:}"

"P_d=\frac{S_d}{S}=\frac{101}{465}"

"\text{probability of delayed company D flight}"

"P_d(D)=\frac{S_d(D)}{S_d}=\frac{59}{101}"

"\text{the probability that the flight is delayed }\newline
\text{and serviced by company D:}"

"P_d*P_d(D)=\frac{101}{465}*\frac{59}{101}\approx0,1269"

Answer:"\approx0,1269"